OpenStudy (anonymous):

Cheyanne plans to put $150 into a savings account. She can place her money into an account represented by f(x) = 5x + 150, or into another account represented by g(x) = 150(1.03)x. Which account has the highest value in 3 years? Which account has the highest value in 10 years?

OpenStudy (anonymous):

(1.03)x the x is an exponent

OpenStudy (anonymous):

@jigglypuff314 @imqwerty @ganeshie8

OpenStudy (bibby):

plug in x=3 and evaluate each function

OpenStudy (anonymous):

f(x) has the highest value in 3 years; g(x) has the highest value in 10 years g(x) has the highest value in 3 years; f(x) has the highest value in 10 years f(x) has the highest value in 3 years; f(x) has the highest value in 10 years g(x) has the highest value in 3 years; g(x) has the highest value in 10 years

OpenStudy (anonymous):

i think its the 3rd one

OpenStudy (bibby):

let's see, it's definitely either a or c because f(x) is at 165 in 3 years

OpenStudy (bibby):

g(x) wins out at 10 years

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

thanks

OpenStudy (bibby):

mhm

OpenStudy (anonymous):

can you help on another one

OpenStudy (bibby):

no promises

OpenStudy (anonymous):

Julie is at the store buying groceries. She needs to get a quart of orange juice and some peaches. The quart of orange juice costs $2.25 and the peaches are $0.75 each. The function f(x) = 0.75x + 2.25 represents Julie's expenses at the grocery store. What do the f(x) and the x represent?

OpenStudy (anonymous):

f(x) represents the number of items purchased at the grocery store; x represents the number of peaches purchased f(x) represents the total cost of the grocery bill; x represents the cost per peach f(x) represents the number of items purchased at the grocery store; x represents the cost per peach f(x) represents the total cost of the groceries; x represents the number of peaches purchased

OpenStudy (anonymous):

those are the options i got d for it is it right?

OpenStudy (bibby):

f(x) represents the total cost of the groceries; x represents the number of peaches purchased

OpenStudy (anonymous):

so i was right :)

OpenStudy (anonymous):

Which of the following exponential functions goes through the points (1, 6) and (2, 12)? f(x) = 3(2)x f(x) = 2(3)x f(x) = 3(2)−x f(x) = 2(3)−x

OpenStudy (anonymous):

all of the last x's on this are exponents

OpenStudy (anonymous):

i got a on this one

OpenStudy (bibby):

looks good, good job

OpenStudy (anonymous):

Marco is studying a type of mold that grows at a fast rate. He created the function f(x) = 345(1.30)x to model the number of mold spores per week. What does the 1.30 represent? How many mold spores are there after 4 weeks? Round your answer to the nearest whole number.

OpenStudy (anonymous):

(1.30)x x is an exponent

OpenStudy (anonymous):

1.30 represents the number of mold spores Marco had at the beginning of the study; 985 mold spores 1.30 represents the number of mold spores grown in a given week; 448 mold spores 1.30 represents the 30% rate of weekly growth; 345 mold spores 1.30 represents the 30% rate of weekly growth; 985 mold spores

OpenStudy (anonymous):

those are the options and i got c

OpenStudy (bibby):

1.3^4=2.8561 345*2.8561=985 spores I think it's d, not positive

OpenStudy (anonymous):

oh ok thank you so much

OpenStudy (anonymous):

that was the last one :)

OpenStudy (bibby):

yay

OpenStudy (anonymous):

thanks i got a 100%